# The Reasoning Bottleneck in Graph-RAG: Structured Prompting and Context Compression for Multi-Hop QA

### 14 Mar 2026 v1

@@ Abstract
Yasaman Zarinkia University of Victoria yasamanz@uvic.ca          Venkatesh Srinivasan Santa Clara University vsrinivasan4@scu.edu          Alex Thomo University of Victoria thomo@uvic.ca
@@ 1 Introduction
Strong retrieval does not guarantee strong answers. Using KET-RAG Huang et al. (2025), a leading Graph-RAG system, with budget open-weight models (Llama-3.1-8B at \$0.05/M tokens, free 384-dim embeddings, total indexing cost $\sim$}\$1 per dataset), we achieve 77% to 91% context coverage across three benchmarks (Table 2), matching the coverage that KET-RAG reports with commercial models (GPT-4o-mini, OpenAI embeddings). Yet overall accuracy ranges from only 23% to 67% for Llama-3.1-8B and 35% to 78% for Llama-3.3-70B. The culprit is context volume: KET-RAG’s retrieved context typically spans $\sim$}10,000 tokens, containing entity descriptions, relationship descriptions, community report summaries, and raw text chunks. The model must locate the relevant facts and chain them together within this large context, a needle-in-a-haystack challenge that intensifies with the number of reasoning hops required. Decomposing errors into retrieval failures (answer absent from context) and reasoning failures (answer present but model incorrect) reveals that reasoning failures account for 73% to 84% of all errors across datasets. This raises a natural question: can inference-time augmentations to the reasoning stage, requiring no retraining or re-indexing, close this gap?
@@
Our augmentations apply to any Graph-RAG system whose retrieved context preserves entity-relationship structure. We focus on KET-RAG and verify that the strategies transfer to LightRAG Guo et al. (2024), another prominent open-source Graph-RAG system. To address this bottleneck, we propose two synergistic strategies that require no additional model training:
@@ 8 Conclusion
Once retrieval achieves high coverage, the dominant bottleneck in Graph-RAG QA is reasoning. We propose three inference-time augmentations that target this gap: SPARQL CoT (+2 to +14 pp), graph-walk compression, and question-type routing; none requires retraining or re-indexing. Together, they enable a fully augmented 8B model to match or exceed an unaugmented 70B baseline on all three benchmarks at ${\sim}12\times$ lower cost, and a replication on LightRAG confirms cross-system transfer. Code and data are available at https://github.com/thomouvic/graph-rag-qa-pub.
@@
Our study has several limitations. First, we evaluate with a single embedding model and two LLM sizes; a finer-grained sweep across model scales and embedding dimensions would better characterize the degradation curve. Second, our context coverage metric uses substring matching, which may overcount (incidental matches) or undercount (paraphrased answers). Third, we use the default $\theta=0.5$ for KET-RAG’s entity-keyword balance, while tuning $\theta$ per dataset might improve results. Finally, our LLM-judged evaluation, while more flexible than exact match, may introduce evaluation noise; we mitigate this with large sample sizes (N=500N{=}500).

### 18 Mar 2026 v2

@@ Abstract
Yasaman Zarrinkia University of Victoria yasamanz@uvic.ca          Venkatesh Srinivasan Santa Clara University vsrinivasan4@scu.edu          Alex Thomo University of Victoria thomo@uvic.ca
@@ 1 Introduction
Strong retrieval does not guarantee strong answers. Using KET-RAG Huang et al. (2025), a state-of-the-art Graph-RAG system, with budget open-weight models (§4), we achieve 77% to 91% context coverage across three benchmarks (Table 2). Yet overall accuracy ranges from only 23% to 67% for Llama-3.1-8B and 35% to 78% for Llama-3.3-70B. The culprit is context volume: KET-RAG’s retrieved context typically spans $\sim$}10,000 tokens, containing entity descriptions, relationship descriptions, community report summaries, and raw text chunks. The model must locate the relevant facts and chain them together within this large context, a needle-in-a-haystack challenge that intensifies with the number of reasoning hops required. Decomposing errors into retrieval failures (answer absent from context) and reasoning failures (answer present but model incorrect) reveals that reasoning failures account for 73% to 84% of all errors across datasets. This raises a natural question: can inference-time augmentations to the reasoning stage, requiring no retraining or re-indexing, close this gap?
@@
To address this bottleneck, we propose two synergistic strategies that require no additional model training. Our augmentations apply to any Graph-RAG system whose retrieved context preserves entity-relationship structure; we focus on KET-RAG and verify that the strategies transfer to LightRAG Guo et al. (2024), another prominent Graph-RAG system.
@@ 8 Conclusion
Once retrieval achieves high coverage, the dominant bottleneck in Graph-RAG QA is reasoning. We propose three inference-time augmentations that target this gap: SPARQL CoT (+2 to +14 pp), graph-walk compression, and question-type routing; none requires retraining or re-indexing. Together, they enable a fully augmented 8B model to match or exceed an unaugmented 70B baseline on all three benchmarks at ${\sim}12\times$ lower cost. Because the graph index is model-independent, stronger future models can be swapped in without re-indexing. For budget deployments, routing with graph-walk compression matches or exceeds 70B quality at a fraction of the cost; for quality-maximizing deployments, SPARQL CoT alone suffices.
Code and data are available at https://github.com/thomouvic/graph-rag-qa-pub.
@@
Our study has several limitations. First, we evaluate with a single embedding model and two LLM sizes; a finer-grained sweep across model scales and embedding dimensions would better characterize the degradation curve. Second, our context coverage metric uses substring matching, which may overcount (incidental matches) or undercount (paraphrased answers). Third, we use the default $\theta=0.5$ for KET-RAG’s entity-keyword balance, while tuning $\theta$ per dataset might improve results. Finally, our LLM-judged evaluation, while more flexible than exact match, may introduce evaluation noise; we mitigate this with large sample sizes (N=500N{=}500). We used an AI coding assistant (Claude Code) to help with experiment scripting and paper editing; all scientific contributions were made by the authors.